# Probing Loop Quantum Gravity via Kerr Black Hole and EHT Results

### v2

@@ Abstract
Recently, a study on shadow of quantum Schwarzschild black hole appeared in [Ye et al., Phys. Lett. B 851, 138566, (2024)] for a fixed value of Barbero-Immirzi parameter $\gamma$. Following this approach, we considered its rotating counterpart being a deformed Kerr metric in Loop Quantum Gravity and studied its deviation from Kerr black hole for a fixed value of $\gamma$. We proved a theorem describing the location of unstable circular null orbits for all metrics of this kind. The deviation between the shadows of the quantum Kerr and Kerr black holes has also been studied, and parameters are constrained by comparison with the EHT results for M87* and Sgr A* to precisely probe the quantity of deviation due to quantum correction. Lastly, we immersed the deformed Kerr black hole in an inhomogeneous plasma and studied its impact on the shadow size. We found that the unstable null orbits for the quantum black hole are always smaller than the unstable null orbits for Kerr black hole. The quantum correction allows the deformed Kerr black hole to mimic Sgr A* with a higher probability than the Kerr black hole. However, the quantum-corrected Kerr black hole barely mimics M87*. The plasma reduces the size of black hole shadow and the plasma parameter in the case II is more sensitive than that in case I.
February 23, 2025
@@ I Introduction
Recently, Lewandowski et al. [28] investigated gravitational collapse of a dust ball by incorporating quantum effects from LQG, using a LQC framework. They found that the collapse halts when the dust ball’s energy density reaches the Planck scale, causing the dust to bounce instead of continuing to collapse. By matching the metrics at the boundary between the collapsing dust ball’s interior and exterior, they derived a LQG-corrected vacuum metric for the external spacetime. During the collapse phase, a quantum-correced Schwarzschild metric appeared, while in the later bouncing phase, features resembling a white hole emerged. This work highlights how quantum gravity influences the behavior of collapsing structures. A few months after this publication, Ye et al. [29] studied the shadow and photon rings of this quantum-corrected BH. He found that the LQG BH can be distinguished from the Schwarzschild BH in terms of shadow images in some illumination models. They assumed the fixed value of the Barbero-Immirzi parameter and hence the parameter $\alpha$ was not treated as a free parameter. Their findings were presented in comparison with the results of Schwarzschild BH. Motivated by this, assuming the fixed value of the Barbero-Immirzi parameter, we will accomplished our analysis in comparison with the results of Kerr BH for the rotating case.
@@
Inspired by this kind of analysis, we consider the recently developed rotating counterpart [77, 78] of the quantum-corrected Schwarzschild BH [28], since the supermassive BHs are rotating in nature. Our major goal is to test the LQG effects under the influence of Barbero-Immirzi parameter, by comparing its shadow results with the EHT data and for Kerr BH. In particular, we investigate the deviation of the quantum-corrected Kerr BH from the classic Kerr BH via EHT results. We also study the impact of plasma on the shadow of deformed Kerr BH. The paper is organized as: In Sect. II, we present a brief overview of the quantum Oppenheimer-Snyder and Swiss-Cheese models and the obtained static BH metric. Further, we also present the rotating BH metric and discuss its horizon structure. In Sect. III, we employ the dynamical methods for the null geodesics, and effective potential and shadows are studied in comparison with result of Kerr BH. The Sect. IV comprises the constraints on the spin of BH parameters in comparison with EHT data. In Sect. V, we investigate the impact of plasma on shadows. Lastly, in Sect. VI, we present a brief conclusion and future prospects. Note that, we consider $G=\hbar=c=1$ in our calculations, unless otherwise mentioned. Moreover, the terms quantum BH, quantum Kerr BH, quantum-corrected BH, deformed BH, deformed Kerr BH, BH in LQG etc. refer to the same BH.
II The Black Hole Metrics in Loop Quantum Gravity
In this section, we review the basic concepts and the development of the quantum-corrected BH Schwarzschild metric and then we derive its rotating counterpart. In Classical GR, the Oppenheimer-Snyder model [79] describes the collapse of dust matter surrounded by the universe described by the Friedmann-Lemaitre-Robertson-Walker metric. Since, the model describes the Big Bang singularity, therefore, it was proposed to consider a Big Bounce instead of the Big Bang [80]. It is also proposed that quantum gravity can resolve the singularity problem, therefore, a Big Bounce is also supported by LQC [12, 81, 82]. The bouncing model can be described by Ashtekar-Pawlowski-Singh (APS) metric [12] given as
@@ II The Black Hole Metrics in Loop Quantum Gravity
The interior region is described by the APS metric, and the exterior one is depicted by qOS metric. The $\theta$ and $\phi$ coordinates are same for both metrics, whereas, the coordinates $\tau$ and $\tilde{r}$ in the ball are matched on to the coordinates $t$ and $r$ in the exterior region. By reversing these regions, a dust ball $0\leq\tilde{r}\leq\tilde{r}_{0}$ is removed from the quantum universe described by the APS metric, now describing the region $\tilde{r}\geq\tilde{r}_{0}$. The empty bubble left over is then filled by the qOS metric which is known as quantum Swiss-Cheese (qSC) model. In this scenario, the bubble contracts before the universe bounces, and it is found that the horizon forms unless the bubble is of the order of the Planck length.
In both cases, the dust interface $\tilde{r}=\tilde{r}_{0}$ is important region in APS spacetime which is described in the other spacetime with coordinates $(t(\tau),r(\tau),\theta,\phi)$. By smoothly matching the spacetimes on the interface so that the coordinates $(\tau,\tilde{r}_{0},\theta,\phi)\sim(t(\tau),r(\tau),\theta,\phi)$ generate the values of the functions $F(r)$ and $H(r)$. As a result, we get the deformed Schwarzschild metric given by [28]
@@
where, the quantum correction parameter $\alpha=16\sqrt{3}\pi\gamma^{3}l_{\text{p}}^{2}$ causes the deformation in the Schwarzschild metric and $l_{\text{p}}=\sqrt{G\hbar}$ defines the Planck length. It is understood that the Barbero-Immirzi parameter has a fixed value $\gamma\approx 0.2375$ [83, 84] that gives a fixed value of the parameter $\alpha\approx 1.1663$ under the assumption $\hbar=G=1=c$ in natural units.
@@
This implies that the radius of the dust surface $a(\tau)\tilde{r}_{0}\in[r_{\text{b}},\infty)$ and thus $r\geq r_{\text{b}}$. The mass $M$ of the dust ball is now the mass of the quantum-corrected BH that has a minimum value $M_{\text{min}}=\frac{16\gamma\sqrt{\pi\gamma}}{3\sqrt[4]{3}}$. Below this minimum value of $M$, there exist no horizon, however, two horizons exist for $M>M_{\text{min}}$ given as
@@
Recently, in Refs. [85, 86, 77, 78], the authors have not considered the above-mentioned fixed value of the Barbero-Immirzi parameter, but accomplished their analyses by incorporating the variation of $\alpha$. It is now well understood that $\gamma$ has a fixed value, therefore, the minimum mass $M_{\text{min}}$ and the deformation parameter $\alpha$ also have fixed values. Therefore, only $M$ is the free parameter in the BH metric (4). The extra term with $\alpha$ describes only the deviation in the quantum regime from the Schwarzschild metric depending on $M$. Generally, for each particular value of $M$, the deviation between the LQG BH and the Schwarzschild BH is fixed due to the fact that $\alpha$ has a fixed value and cannot be varied around its prescribed value. Therefore, the studies mentioned above with variable $\alpha$ do not account for rigorous and feasible analyses. Since, $M>M_{\text{min}}\approx 0.8314$ is a valid limit for the BH mass, therefore, we consider $M=1$ throughout this work to focus on other parameters.
The significance of studying rotating BHs can be highlighted through a comparison of their shadows with the EHT data on supermassive BHs as they are predominantly rotating in nature. Thus, considering rotating BHs allow us to establish a more accurate and rigorous analysis. The rotating counterpart of the metric (4) can be easily derived by employing a well-known modified Newman-Janis algorithm [87, 88]. The rotating BH metric in Boyer-Lindquist coordinates reads [77, 78]
@@
where, $a$ is the spin parameter. Recently, this solution also appeared in the arXiv submissions [77, 78], however, they have treated the LQG parameter $\alpha$ as a variable or free parameter. We know that a matter source corresponding to a static BH may get altered with some additional sources due to the modified Newman-Janis algorithm. Therefore, the new matter sources corresponding to the rotating BH are often unknown. It was verified by Hansen and Yunes [89] for the case of modified gravity. Consequently, the obtained metric may not be exact rotating metric for its static counterpart. In this context, one may also not be able to find the modification in the source because of the unknown terms in the action and the energy-momentum tensor. In our case, the static metric was derived by considering the collapse and matching conditions. Therefore, the action and field equations being unknown, one cannot assess whether the initially collapsing dust ball remains preserved or is altered with some new sources. For this reason, we have marked the metric (7) as an effective metric which encompasses the most probable properties of the exact rotating metric. For instance, the metric (7) reduces to the metric (4) when $a=0$. By removing the quantum effects, the metric (7) reduces to the Kerr metric, just as the metric (4) reduces to Schwarzschild one. Moreover, the metric function $\Delta(r)$ can be written as
@@
which shows that the metric (7) is a deformed Kerr BH metric with an additional quantum correction term. We have mentioned that the parameter $\alpha$ has a fixed value and we consider $M=1$, thus, the only free parameter for the rotating metric is the spin $a$. Hence, we can rigorously analyze the deviation of quantum-corrected Kerr BH from the Kerr BH. Like the Kerr metric, the deformed Kerr BH metric also exhibits the time-translational and rotational invariance isometries corresponding to the Killing vector fields $\left(\partial_{t}\right)^{\mu}$ and $\left(\partial_{\phi}\right)^{\mu}$.
The last term in Eq. (10) has no dependence on $a$, therefore, the horizon curve will be unique with a certain deviation from the horizon curve of Kerr BH. It is generated numerically by solving the equation $\Delta(r)=0$ for real and positive roots. A clear deviation in two curves is shown in Fig. 1 in terms of horizon radii $r_{\text{h}}$ with respect to $a$. It can be seen that the event horizon of the deformed Kerr BH is smaller than the event horizon of Kerr BH. While the Cauchy horizon of Kerr BH is significantly smaller than the event horizon of quantum Kerr BH. Moreover, the extremal deformed Kerr BH has spin $\sim$0.4952 which is marginally less than half of the extremal spin of Kerr BH.
Figure 1: Horizon structure of Kerr and deformed Kerr BHs with respect to $a$ for $M=1$ and $\alpha\approx 1.1663$.
@@ III Unstable Null Orbits and Black Hole Shadow
The photons emerging from a bright source may get trapped in unstable and stable circular null orbits in the vicinity of a BH. Some of them in the unstable orbits fall into the event horizon, while the rest scatter away to the infinity. This is how the optical image of the BH is formed, termed as shadow [90, 91]. These orbits are characterized by an effective potential function. Therefore, we derive the null geodesic equations to study the effective potential and shadows of the Kerr BH in LQG, in order to understand the influence of LQG on deformed Kerr BH in terms of unstable orbits and optical images. For the Kerr BH in LQG, the null geodesic equations can be obtained by employing the Hamilton-Jacobi formalism [92] which has been widely incorporated in the literature over the years. One can also begin with the Lagrangian or Hamiltonian methods, generating two constants of motion, the energy $E$ and the angular momentum $L$ along with mass of the particle under motion. The geodesic equations become completely integrable if we introduce a fourth constant of motion. The separation of variables arising in Hamilton-Jacobi formalism enables us to introduce a constant known as the Carter constant [92]. We consider the Hamilton-Jacobi equation
@@
Theorem: Suppose a Kerr-like BH described by $\Delta(x)=\Delta_{\text{Kerr}}(x)+a_{1}(x)$ with $V^{\text{eff}}(x)=V_{\text{Kerr}}^{\text{eff}}(x)+a_{2}(x)$ such that $a_{1}(x)=b_{1}x^{p}$ and $a_{2}(x)=b_{2}x^{q}$ be decreasing functions in an open interval $\mathcal{I}=(0,s)\subseteq\mathbb{R}^{+}$, and $b_{1},b_{2}$ be positive real constants. If an unstable null orbit for Kerr BH exists at a radial distance $x_{0}\in\mathcal{I}$, then the unstable null orbit corresponding to $\Delta(x)$ exists at some $y_{0}\in\mathcal{I}$ such that $y_{0}<x_{0}$.
Figure 2: Behavior of effective potential and unstable null orbits for Kerr and deformed Kerr BHs for different values of $a$ with $M=1$ and $\alpha\approx 1.1663$.
Proof: To prove the theorem, we consider the convexity of the effective potential functions. Given that the metric function and the effective potential function for Kerr-like BH can be decomposed into the metric function and the effective potential function for Kerr BH with decreasing functions $a_{1}(x)$ and $a_{2}(x)$, respectively. It implies that the exponents $p,q\in\mathbb{R}^{-}$ and the constants $b_{1},b_{2}$ may not contain any free parameter except for spin parameter $a$. Now, if an unstable null circular orbit for Kerr BH exists at a radial distance $x_{0}\in\mathcal{I}$, then by definition, the function $V_{\text{Kerr}}^{\text{eff}}(x)$ is concave at $x_{0}$ which implies that $\partial_{x}^{2}V_{\text{Kerr}}^{\text{eff}}(x_{0})<0$ and $\partial_{x}V_{\text{Kerr}}^{\text{eff}}(x_{0})=0$. Since, $a_{2}(x)$ is decreasing $\forall~{}x\in\mathcal{I}$, therefore, $\partial_{x}V^{\text{eff}}(x_{0})<0$. From this, one of the following two possibilities must hold:
@@
    If $V^{\text{eff}}(x)$ is concave at $y_{0}$, then $\exists~{}y_{0}\in\mathcal{I}$ such that $y_{0}<x_{0}$ and $\partial_{x}V^{\text{eff}}(y_{0})=0$.
@@
    If $V^{\text{eff}}(x)$ is convex at $y_{0}$, then $\exists~{}y_{0}\in\mathcal{I}$ such that $y_{0}>x_{0}$ and $\partial_{x}V^{\text{eff}}(y_{0})=0$.
We need to prove $(i)$ to prove the theorem or else we must show that $(ii)$ does not hold. Suppose that $(ii)$ holds, then $\partial_{x}V_{\text{Kerr}}^{\text{eff}}(y_{0})>0$ but $V_{\text{Kerr}}^{\text{eff}}(x)$ must be decreasing at $y_{0}$ under the given statement. This gives a contradiction which implies that $(ii)$ does not hold and thus $(i)$ holds automatically. This proves the theorem that the unstable null orbits shrink under the given assumptions. One may point out that we have not proved the existence of $y_{0}$ in $\mathcal{I}$. Since, the lower bound of the interval $\mathcal{I}$ is 00 and $y_{0}<x_{0}$, therefore, $y_{0}\in\mathcal{I}$. We have also omitted the case when $y_{0}=x_{0}$ because $b_{1},b_{2}\neq 0$. Moreover, there is no restriction on the interval $\mathcal{I}$ as it may have multiple extrema, however, the local maximum corresponding to the unstable orbit is unique.
The result proved in the above theorem can be viewed in Fig. 2 in terms of location of peaks of the curves. That is, the unstable null orbits for quantum Kerr BH are smaller than the unstable null orbits for Kerr BH for all cases. Moreover, for each case of spin up to extremal quantum Kerr BH, the difference between the sizes of unstable orbits for Kerr and quantum Kerr BHs increases, which is specifically due to the presence of spin parameter in the deviation term in Eq. (20). Furthermore, with increase in spin, the unstable orbits of quantum and Kerr BHs shrink.
@@
Figure 3: Behavior of shadows for Kerr and deformed Kerr BHs for different values of $a$ with $M=1$ and $\alpha\approx 1.1663$, visualized by an equatorial observer at radial infinity.
@@
which further reduce to the celestial coordinates for Kerr BH under the limit $\alpha=0$. The extra terms on the right hand side of Eqs. (27) and (28) describe the deviation of shadow due to LQG effects from the shadow of Kerr BH. The $Y$-coordinate in Eq. (28) is expressed in terms of an alternating infinite series contributing to the deviation of the shadow from the shadow of Kerr BH. To quantify this deviation in the shadow, we plot some shadow contours for a few cases of spin parameter values as given in Fig. 3. From the difference of sizes of the quantum Kerr and Kerr BHs for all cases, one can deduce that the quantum Kerr BH will appear smaller to a visualizing observer at infinity as compared to the Kerr BH. As the value of spin parameter increases, the distortion starts appearing in the shadows of quantum Kerr BH. That is, the difference between the shadow contours on the left side increases with increase in spin. It suggests that the spin parameter influences the elongation of shadows for quantum Kerr BH more than the Kerr BH up to $a=0.4952$. The third plot with three contours show a huge difference between the shadows of quantum and Kerr BHs for their respective extreme spin values. Therefore, one may not expect to visualize a perfectly flattened shadow for extreme LQG BH. Moreover, the deviation in shadow contours for $a=0.4952$ for both BHs can also be measured. For smaller values of spin, both BHs are centered closer to the origin and will certainly be centered at origin for $a=0$.
Figure 4: Comparison of shadow angular diameter $\theta_{\text{d}}$ for quantum Kerr BH (dashed red curve) and Kerr BH (solid black curves) with the EHT data for M87* (at inclination angle of $17^{\circ}$) and Sgr A* (at inclination angle of $45^{\circ}$) for the bounds on spin $a$ within 1-$\sigma$ intervals.

### v3

# Influence of Quantum Correction on Kerr Black Hole in Effective Loop Quantum Gravity via Shadows and EHT Results

@@ Abstract
Recently, a study on shadow of quantum corrected Schwarzschild black hole in loop quantum gravity appeared in [Ye et al., Phys. Lett. B 851, 138566, (2024)] assuming a fixed value of Barbero-Immirzi parameter $\gamma$. Following this approach, we considered its rotating counterpart being a quantum corrected Kerr black hole in effective loop quantum gravity and studied its deviation from Kerr black hole for a fixed value of $\gamma$. We proposed and proved a theorem describing the location of unstable circular null orbits for all such Kerr-like metrics. The deviation between the shadows of the Kerr and quantum corrected Kerr black holes has also been studied, and parameters are constrained by comparison with the EHT results for M87* and Sgr A* to precisely probe the quantity of deviation due to quantum correction. Lastly, we immersed the quantum corrected Kerr black hole in an inhomogeneous plasma and studied its impact on the shadow size. We found that the unstable null orbits for the quantum corrected Kerr black hole are always smaller than the unstable null orbits for Kerr black hole. The effect of Barbero-Immirzi parameter allows the quantum corrected Kerr black hole to mimic Sgr A* with a higher probability than the Kerr black hole. However, the quantum corrected Kerr black hole does not mimic M87*. The plasma reduces the size of the shadow of quantum corrected black hole, and the plasma parameter in the case II is more sensitive than that in case I.
March 14, 2025
@@ I Introduction
Recently, Lewandowski et al. [28] investigated gravitational collapse of a dust ball by incorporating quantum effects from LQG, using a LQC framework. They found that the collapse halts when the dust ball’s energy density reaches the Planck scale, causing the dust to bounce instead of continuing to collapse. By matching the metrics at the boundary between the collapsing dust ball’s interior and exterior, they derived a quantum correced Schwarzschild metric for the external spacetime. This work highlights how quantum gravity influences the behavior of collapsing structures. Later, Ye et al. [29] studied the shadow and photon rings of this quantum corrected BH. He found that this quantum corrected BH can be distinguished from the Schwarzschild BH in terms of shadow images in some illumination models. They assumed the fixed value of the Barbero-Immirzi parameter and hence the parameter $\alpha$ was not treated as a free parameter. Their findings were presented in comparison with the results of Schwarzschild BH. Motivated by this, assuming the fixed value of the Barbero-Immirzi parameter, we will accomplish our analysis in comparison with the results of Kerr BH for the rotating case.
@@
Inspired by [29], we consider the recently developed rotating counterpart [77, 78, 79] of the quantum corrected Schwarzschild BH [28], since the supermassive BHs are rotating in nature. Our major goal is to test the LQG effects under the influence of Barbero-Immirzi parameter, by comparing its shadow results with the EHT data and for Kerr BH. In particular, we investigate the deviation of the quantum corrected Kerr BH from the Kerr BH via EHT results. We also study the impact of plasma on the shadow of quantum corrected Kerr BH. The paper is organized as: In Sect. II, we present a brief overview of the static and rotating BH metrics, and further discuss its horizon structure in terms of BH spin. In Sect. III, we employ the dynamical methods for the null geodesics, and effective potential and shadows are studied in comparison with result of Kerr BH. The Sect. IV comprises the constraints on the spin of BH parameters in comparison with EHT data. In Sect. V, we investigate the impact of plasma on shadows. Lastly, in Sect. VI, we present a brief conclusion and future prospects. Note that, we consider $G=\hbar=c=1$ in our calculations, unless otherwise mentioned.
II The Static and Rotating Quantum Corrected Black Hole Metrics
In this section, we review the basic concepts and the development of the quantum corrected Schwarzschild BH metric and its rotating counterpart. The Oppenheimer-Snyder model [80] describes the collapse of dust matter. Since, the model describes the Big Bang singularity, therefore, it was proposed to consider a Big Bounce instead of the Big Bang [81]. It is also proposed that quantum gravity can resolve the singularity problem, therefore, a Big Bounce is also supported by LQC [12, 82, 83]. A 4D spherically symmetric Ashtekar-Pawlowski-Singh (APS) metric [12] given as
@@ II The Static and Rotating Quantum Corrected Black Hole Metrics
The interior region is described by the APS metric as a dust ball, and the exterior one is depicted by vacuum qOS metric. The $\theta$ and $\phi$ coordinates are same for both metrics, whereas, the coordinates $\tau$ and $\tilde{r}$ in the ball are matched on to the coordinates $t$ and $r$ in the exterior region. The dust interface $\tilde{r}=\tilde{r}_{0}$ is important region in APS spacetime, with coordinates $(t(\tau),r(\tau),\theta,\phi)$ in the other spacetime. Using the Israel junction conditions, which require continuity of the metrics and extrinsic curvature on the interface between dust and vacuum, the coordinates are matched as $(\tau,\tilde{r}_{0},\theta,\phi)\sim(t(\tau),r(\tau),\theta,\phi)$, generating a quantum corrected Schwarzschild metric in LQG give by [28]
@@
where, the quantum correction parameter $\alpha=16\sqrt{3}\pi\gamma^{3}l_{\text{p}}^{2}$ causes the deformation in the Schwarzschild metric, and $l_{\text{p}}=\sqrt{G\hbar}$ defines the Planck length. It is understood that the Barbero-Immirzi parameter has a fixed value $\gamma\approx 0.2375$ [84, 85] that gives a fixed value of the parameter $\alpha\approx 1.1663$ under the assumption $\hbar=G=1=c$ in natural units.
@@
This implies that the radius of the dust surface $a(\tau)\tilde{r}_{0}\in[r_{\text{b}},\infty)$ and thus $r\geq r_{\text{b}}$. The mass $M$ of the dust ball is now the mass of the quantum corrected BH that has a minimum value $M_{\text{min}}=\frac{16\gamma\sqrt{\pi\gamma}}{3\sqrt[4]{3}}$. Below this minimum value of $M$, there exist no horizon, however, two horizons exist for $M>M_{\text{min}}$ given as
@@
Recently, in Refs. [86, 87, 78, 79], the authors have not considered the above-mentioned fixed value of the Barbero-Immirzi parameter, but accomplished their analyses by incorporating the variation of $\alpha$. It is now well understood that $\gamma$ has a fixed value, therefore, the minimum mass $M_{\text{min}}$ and the deformation parameter $\alpha$ also have fixed values. Therefore, only $M$ is the free parameter in the BH metric (4). The extra term with $\alpha$ describes only the deviation in the quantum regime from the Schwarzschild metric depending on $M$. Generally, for each particular value of $M$, the deviation between the quantum corrected BH and the Schwarzschild BH is fixed due to the fact that $\alpha$ has a fixed value and cannot be varied around its prescribed value. Therefore, the studies mentioned above with variable $\alpha$ do not account for rigorous and feasible analyses. Since, $M>M_{\text{min}}\approx 0.8314$ is a valid limit for the BH mass, therefore, we consider $M=1$ throughout this work to focus on influence of $a$ within the LQG regime.
The significance of studying rotating BHs can be highlighted through a comparison of their shadows with the EHT data on supermassive BHs as they are predominantly rotating in nature. Thus, considering rotating BHs allow us to establish a more accurate and rigorous analysis. The rotating counterpart of the metric (4) with effective geometry was derived in Refs. [77, 78, 79] so that the quantum corrected Kerr BH in effective LQG in Boyer-Lindquist coordinates reads
@@
where, $a$ is the spin parameter. In the static case, the model does not admit classical covariance. However, it is proved that there is a modified form of covariance, since the constraint algebra closes, even if the structure functions are modified [88]. Since, the covariance is modified in static case for the spherical symmetry, the rotating metric may not be exact line element for its static counterpart giving rise to an effective geometry. Moreover, it is not required at all that classical covariance needs to be preserved at the effective level. For this reason, the metric (7) serves as an effective metric and therefore describes a quantum corrected Kerr BH in effective LQG which encompasses the most probable properties of the exact rotating metric. For instance, the metric (7) reduces to the metric (4) when $a=0$. By removing the quantum effects, the metric (7) reduces to the Kerr metric, just as the metric (4) reduces to Schwarzschild one. Moreover, the metric function $\Delta(r)$ can be written as
@@
which shows that the metric (7) is a quantum corrected Kerr BH metric with an additional term comprising LQG parameter. We have mentioned that the parameter $\alpha$ has a fixed value and we consider $M=1$, thus, the only free parameter for the rotating metric is the spin $a$. Hence, we can rigorously analyze the deviation of quantum corrected Kerr BH from the Kerr BH. Like the Kerr metric, the quantum corrected Kerr metric also exhibits the time-translational and rotational invariance isometries corresponding to the Killing vector fields $\left(\partial_{t}\right)^{\mu}$ and $\left(\partial_{\phi}\right)^{\mu}$.
The last term in Eq. (10) has no dependence on $a$, therefore, the horizon curve will be unique with a certain deviation from the horizon curve of Kerr BH. It is generated numerically by solving the equation $\Delta(r)=0$ for real and positive roots. A clear deviation in two curves is shown in Fig. 1 in terms of horizon radii $r_{\text{h}}$ with respect to $a$. It can be seen that the event horizon of the quantum corrected Kerr BH is smaller than the event horizon of Kerr BH. While the Cauchy horizon of Kerr BH is significantly smaller than the event horizon of quantum corrected Kerr BH. Moreover, the extremal quantum corrected Kerr BH has spin $\sim$0.4952 which is marginally less than half of the extremal spin of Kerr BH.
Figure 1: Horizon structure of Kerr and quantum corrected Kerr BHs with respect to $a$ for $M=1$ and $\alpha\approx 1.1663$.
@@ III Unstable Null Orbits and Black Hole Shadow
The photons emerging from a bright source may get trapped in unstable and stable circular null orbits in the vicinity of a BH. Some of them in the unstable orbits fall into the event horizon, while the rest scatter away to the infinity. This is how the optical image of the BH is formed, termed as shadow [89, 90]. These orbits are characterized by an effective potential function. Therefore, we derive the null geodesic equations to study the effective potential and shadows of quantum corrected Kerr BH in effective LQG, in order to understand the influence of LQG via Barbero-Immirzi parameter in terms of unstable orbits and optical images. For the quantum corrected Kerr BH in effective LQG, the null geodesic equations can be obtained by employing the Hamilton-Jacobi formalism [91] which has been widely incorporated in the literature over the years. One can also begin with the Lagrangian or Hamiltonian methods, generating two constants of motion, the energy $E$ and the angular momentum $L$ along with mass of the particle under motion. The geodesic equations become completely integrable if we introduce a fourth constant of motion. The separation of variables arising in Hamilton-Jacobi formalism enables us to introduce a constant known as the Carter constant [91]. We consider the Hamilton-Jacobi equation
@@
Theorem: Suppose a Kerr-like metric described by $\Delta(r)=\Delta_{\text{Kerr}}(r)+b_{1}r^{p}$ with $V^{\text{eff}}(r)=V_{\text{Kerr}}^{\text{eff}}(r)+b_{2}r^{q}$ such that $r^{p}$ and $r^{q}$ be decreasing functions in an open interval $\mathcal{I}=(0,s)\subseteq\mathbb{R}^{+}$, and $b_{1},b_{2}\in\mathbb{R}^{+}$. If an unstable null orbit for Kerr BH exists at a radial distance $r_{k}\in\mathcal{I}$, then the unstable null orbit corresponding to $\Delta(r)$ exists at some $r_{0}\in\mathcal{I}$ such that $r_{0}<r_{k}$.
Figure 2: Behavior of effective potential and unstable null orbits for Kerr and quantum corrected Kerr BHs for different values of $a$ with $M=1$ and $\alpha\approx 1.1663$.
Proof: Given that $\Delta(r)=\Delta_{\text{Kerr}}(r)+b_{1}r^{p}$ and $V^{\text{eff}}(r)=V_{\text{Kerr}}^{\text{eff}}(r)+b_{2}r^{q}$ with decreasing functions $r^{p}$ and $r^{q}$ such that $p,q\in\mathbb{R}^{-}$, and $b_{1},b_{2}\in\mathbb{R}^{+}$ depend only on spin $a$. Now, if an unstable null orbit for Kerr BH exists at some $r_{k}\in\mathcal{I}$, then the function $V_{\text{Kerr}}^{\text{eff}}(r)$ is concave at $r_{k}$ that is $\partial_{r}^{2}V_{\text{Kerr}}^{\text{eff}}(r_{k})<0$ and $\partial_{r}V_{\text{Kerr}}^{\text{eff}}(r_{k})=0$. Since, $r^{q}$ is decreasing $\forall~{}r\in\mathcal{I}$, therefore, $\partial_{r}V^{\text{eff}}(r_{k})<0$. From this, one of the following two possibilities holds:
@@
    If $V^{\text{eff}}(r)$ is concave at $r_{0}$, then $\exists~{}r_{0}\in\mathcal{I}$ such that $r_{0}<r_{k}$ and $\partial_{r}V^{\text{eff}}(r_{0})=0$.
@@
    If $V^{\text{eff}}(r)$ is convex at $r_{0}$, then $\exists~{}r_{0}\in\mathcal{I}$ such that $r_{0}>r_{k}$ and $\partial_{r}V^{\text{eff}}(r_{0})=0$.
To prove the theorem, we either prove $(i)$ holds or $(ii)$ does not hold. Contrarily suppose that $(ii)$ holds, then $\partial_{r}V_{\text{Kerr}}^{\text{eff}}(r_{0})>0$ but $V_{\text{Kerr}}^{\text{eff}}(r)$ must be decreasing at $r_{0}$ under the given statement. This gives a contradiction which implies that $(ii)$ does not hold and thus $(i)$ holds automatically. Therefore, it proves that the unstable null orbits shrink under the given assumptions. Note that we have not proved $r_{0}\in\mathcal{I}$. Since, the lower bound of the interval $\mathcal{I}$ is 00 and $r_{0}<r_{k}$, therefore, $r_{0}\in\mathcal{I}$. We have also omitted the case when $r_{0}=r_{k}$ because $b_{1},b_{2}\neq 0$. Moreover, the interval $\mathcal{I}$ is arbitrary and may have multiple extrema, however, the local maximum corresponding to the unstable orbit is unique.
The result proved in the above theorem can be viewed in Fig. 2 in terms of location of peaks of the curves. That is, the unstable null orbits for quantum corrected Kerr BH are smaller than the unstable null orbits for Kerr BH for all cases. Moreover, for each case of extremal spin of quantum corrected Kerr BH, the difference between the sizes of unstable orbits for both BHs increases, which is specifically due to the presence of spin parameter in the deviation term in Eq. (20). Furthermore, with increase in spin, the unstable orbits of quantum corrected and Kerr BHs shrink.
@@
Figure 3: Behavior of shadows for Kerr and quantum corrected Kerr BHs for different values of $a$ with $M=1$ and $\alpha\approx 1.1663$, visualized by an equatorial observer at radial infinity.
@@
which further reduce to the celestial coordinates for Kerr BH under the limit $\alpha=0$. The extra terms on the right hand side of Eqs. (27) and (28) describe the deviation of shadow due to LQG effects from the shadow of Kerr BH. The $Y$-coordinate in Eq. (28) is expressed in terms of an alternating infinite series contributing to the deviation of the shadow from the shadow of Kerr BH. To quantify this deviation in the shadow, we plot some shadow contours for a few cases of spin parameter values as given in Fig. 3. From the difference of sizes of the Kerr and quantum corrected Kerr BHs for all cases, one can deduce that the quantum corrected Kerr BH will appear smaller to a visualizing observer at infinity as compared to the Kerr BH. As the value of spin parameter increases, the distortion starts appearing in the shadows of quantum corrected Kerr BH. That is, the difference between the shadow contours on the left side increases with increase in spin. It suggests that the spin parameter influences the elongation of shadows for quantum corrected Kerr BH more than the Kerr BH up to $a=0.4952$. The third plot with three contours show a huge difference between the shadows of quantum corrected and Kerr BHs for their respective extreme spin values. Therefore, one may not expect to visualize a perfectly flattened shadow for extreme quantum corrected Kerr BH. Moreover, the deviation in shadow contours for $a=0.4952$ for both BHs can also be measured. For smaller values of spin, both BHs are centered closer to the origin and will certainly be centered at origin for $a=0$.
Figure 4: Comparison of shadow angular diameter $\theta_{\text{d}}$ for quantum corrected Kerr BH (dashed red curve) and Kerr BH (solid black curves) with the EHT data for M87* (at inclination angle of $17^{\circ}$) and Sgr A* (at inclination angle of $45^{\circ}$) for the bounds on spin $a$ within 1-$\sigma$ intervals.